# Building an Open AIBOM Standard in the Wild

### v1

@@ 5.1. Regulatory and Standards Alignment
Methodology. We analyzed the alignment between fields captured in the AIBOM specification and the regulatory and ethical requirements defined by the standards and regulations shown in Table 2. The analysis was performed by the fourth and fifth author, who have extensive experience in standards development and regulatory alignment, and was further independently reviewed by first and the sixth authors for additional validation.
EU AI Act., we examined the obligations defined in Articles 49 and 60, which govern mandatory registration in the EU high-risk AI database prior to deployment or testing. Required information was grouped into four categories: Identification, System Details, Verification Details, and Application Details, comprising a total of 14 subcategories (e.g., provider contact, intended purpose, classification, testing plan, and involved parties).
US and EU medical device regulations, we analyzed our AIBOM specification’s alignment with Article 10 of Regulation (EU) 2017/745 and FDA cybersecurity guidance. Information elements were organized into three categories: package details, model details, and data details.
@@
Results. Our analysis shows that the SPDX 3.0 AIBOM specification provides strong coverage of regulatory, safety, and ethical requirements. We found that we could successfully represent 13 of 14 information obligations under the EU AI Act, capturing all required elements from US and EU medical device regulations. Our validation shows that AIBOM specification aligns with over 40 subclauses across eight IEEE 7000 series standards i.e., more than 90% overall coverage indicating practical readiness for compliance adoption.
@@ 5.2. Validation against Key Industry Use Cases
Motivation. While regulatory alignment ensures that our AIBOM specification meets legal and compliance obligations, its practical value ultimately depends on how well it addresses real-world needs identified by industry stakeholders. In June 2025, the SBOM for AI Tiger Team (SBOM for AI Tiger Team, 2025), convened by the U.S. Cybersecurity and Infrastructure Security Agency (CISA) [^2] https://www.cisa.gov/, accessed September 30, 2025, articulated six foundational use cases that any AI-focused SBOM standard must support. These use cases span the full lifecycle of AI system deployment: regulatory compliance, Vulnerability and Incident Management, legal and intellectual property assurance, third-Party AI Risk Management, Open Source Model Risk Assessment and Model Lifecycle and Asset Management. A detailed overview of these usecases can be found here (Use, 2025).
Validating our AIBOM specification against these use cases is therefore essential to demonstrate that the specification is not only standards-compliant but also operationally relevant, usable by organizations in practice, and capable of supporting the core security, legal, and governance workflows envisioned by the broader AIBOM consumer ecosystem.
Methodology. The CISA “SBOM for AI” use cases are intentionally high-level; however, assessing whether our AIBOM specification is capable of capturing the metadata outlined in them requires concrete, testable requirements. We therefore operationalized these use cases by consolidating them into broader themes, assembling a focused evidence base from recent studies, extracting atomic requirements, and mapping them to SPDX 3.0 AIBOM fields before performing external validation.
@@
Step 2: Build an evidence base per theme. For each theme, we identified relevant research papers and technical reports that analyze concrete instances of the use cases. For example, Rajbahadur et al. (Rajbahadur et al., 2021), which outlines methods for conducting dataset license compliance analysis, was mapped to the Compliance category. All selected studies are listed in Table 3.
@@
Step 5: Internal and external validation. The first author independently verified the extractions and mappings. We then presented the traceability matrix and representative examples to the AI SBOM Tiger Team (29 Sept 2025, attended by 12 participants) and the SPDX AI and Dataset WG (attended by four participants). Authors of this paper who were part of the working group did not participate in the validation exercise. Feedback was solicited on whether AIBOM could be effectively operationalized to address the six use cases. Meeting minutes from both working group sessions are publicly available. [^3] SBOM for AI (AIBOM) Tiger Team working group meeting minutes https://docs.google.com/document/d/1IpXG7XBOJnPl_hwFf3JZkDaFb0k2CnI0/edit?usp=sharing&ouid=110194678381965933391&rtpof=true&sd=true
Results. Our analysis demonstrated that the AIBOM specification could represent all 46 distinct requirements extracted from the eight studies reviewed across the three consolidated use-case categories. Table 3 summarizes the number of AIBOM fields across the AI, Dataset, and other SPDX profiles that satisfied the requirements identified in each study. In several cases, a single requirement could be addressed by multiple fields, underscoring the flexibility and composability of our AIBOM specification. A detailed traceability matrix, including the full set of extracted requirements and their field-level mappings, is available in our extended technical report (Rashno, 2025).
@@
Open Source Risk Management (Use, 2023b) 6 5 4
@@ 5.3. Validation based on Practitioner Interviews
Methodology. It is important to note that this validation was conducted prior to the official release of the AIBOM standard. This timing ensured that participants focused on the fields they considered essential for an effective AIBOM, rather than merely commenting on the limitations of our finalized specification. A detailed overview of our methodology is presented below.
Participant Recruitment. For this study, we recruited practitioners with experience working with SBOM or developing software with an AI component. We relied on our professional networks to recruit participants for this study. Throughout the 2022-mid 2023 period, we directly emailed experts who are known to have experience with SBOMs or developing AI software. From the individuals with relevant experience who responded to our emails, we ultimately selected ten participants for our study.
Table 4 shows the demographic details of the ten participants. The interview participants varied in their work experience, work domain, and their future relationship with AIBOM.
@@
Interview Protocol. First, we drafted an interview questionnaire and presented it to two experts from academia and industry. Based on their feedback, we finalized the interview questions. The 18 questions in the interviews centered around the following themes: the participants’ past experience with AI and software development, a working definition of AI software/system traceability, current practices and tools available for ensuring AI system traceability, and expectations from a new AIBOM specification.
We conducted semi-structured interviews with each participant as it allowed for improvisation and exploration of the studied objects (Wohlin et al., 2012). All the interviews were conducted via video calls and lasted 30 to 60 minutes each. To avoid bias, we did not show the proposed fields to the participants. We received consent from all participants to record the interviews and use the recordings for research purposes. We transcribed the recordings using software and manually checked for errors.
Transcript Coding and Analysis. First, two authors sat together to read through two interview transcripts in full and induce codes from interview passages. Then, the authors independently read the remaining interview transcripts thoroughly and assigned codes to passages. Next, the authors met again to filter out the codes irrelevant to our goal and grouped the independently identified codes for each transcript. Once a consensus was reached about the codes and categories from the interviews, we computed the overlap between the fields included in our AIBOM specification and the concepts proposed by the interview participants.
@@
Results. With the exception of Configuration and Deployment-related fields (which we had deliberately earmarked for inclusion in a future version due to difficulty with capturing them) all fields proposed by participants could be represented within the current SPDX AIBOM 3.0 specification. Table 5 presents the mapping between the fields suggested by study participants and those included in the AIBOM specification.
The primary theme emerging from the interviews was the widespread inadequacy of current traceability practices for AI systems, which participants described as ad hoc and insufficient for ensuring proper governance. This sentiment was powerfully articulated by an executive director (P2), who stated, ”I think it’s a general mystery across the industry… I don’t think that [traceability] adequately exists within the AI community around algorithms and training sets to be able to demonstrate that traceability from an auditing standpoint.” This lack of a standardized, auditable record is exacerbated by real-world development pressures. A machine learning engineer (P7) from a startup described this vividly: ”It’s something that we lack and it’s because… the general nature of a startup, we usually deprioritize documentation and it usually also ends with a lot of pain for us… because we don’t have the documentation, I usually spend so much time trying to explain anyway.”
Beyond identifying these gaps, participants also provided clear insights into what an effective traceability solution should capture. When asked what information they need when consuming a pre-trained model, a lead data scientist (P1) offered a detailed wishlist: ”What do I look for? I look for license… support… When, which training data was used, what demographic was used, and what biases do they have? …And these are stuff that you want captured in the document, because that’s what I’m looking for [as a model consumer]… what’s the [reported] accuracy? How did you test it?” This practitioner’s desired list of artifacts, spanning licensing, data provenance, bias, and evaluation metrics, directly aligns with the core categories of the proposed AIBOM framework, confirming that its design addresses the expressed needs of practitioners. These insights will continue to guide the future evolution of our AIBOM specification.
@@ 5.4. Field Validation in Industrial Setting
We report an external study conducted at a large multinational software organization to evaluate AIBOM’s suitability for representing AI system documentation use cases in practice. The study was conducted by a team of security architects, data scientists, DevOps experts, and AI developers. From June 2024 to September 2024, the team members worked on identifying the overlap between the fields available in our AIBOM specification and the fields encountered by practitioners during the organization’s development of an AI system.
Methodology. The organization already had a checklist for developers to evaluate how their AI software development affects the internal ethics policy. Moreover, the legal department had a checklist to assess the legal impact of each AI-related software release. The researchers investigated how many of the fields in these checklists could be populated if the organizations already had AIBOMs constructed as part of the AI software development process. Furthermore, they investigated whether our AIBOM specification can be used to automatically generate model cards. Then, they investigated to what extent AIBOM generation for third-party models can be automated using web scraping or external APIs, such as Hugging Face (Hugging Face, 2025).
Results. The analysis found that all of the fields required in internal checklists in the organization’s AI system development process can be populated from the AIBOM specification’s fields, demonstrating the comprehensiveness of AIBOM in practice. They observed that 60% of the fields in the existing model cards could be extracted using the fields in the AIBOM specification. Furthermore, they found that external APIs and web scraping could populate 40% of the fields when generating AIBOMs for third-party models used at the organization. This demonstrated that AIBOM generation could be partially automated for practical use cases, reducing the manual effort needed.
@@ 6. Standards Development as Action Research
A key contribution of our paper is demonstrating how the development of a complex, multi-stakeholder standard can be systematically understood through the lens of Action Research (AR). AR is a methodology that addresses real-world problems while producing actionable knowledge (Baskerville, 1999; Elden and Chisholm, 1993; Ferrario et al., 2014). By retrospectively framing our large-scale effort to extend SPDX and create an AIBOM specification as an AR process, we provide a concrete example of how community-driven standards development can follow a disciplined, research-grounded approach in rapidly evolving domains.
Although we did not initially conceive this project as AR, its conduct and outcomes closely align with AR principles. AR is fundamentally concerned with changing practice, with an “explicit aim to act in the real world and change the state of practice.” Our goal was to create a functional, community-driven standard for AI transparency and compliance i.e., an intervention by design. We therefore frame our work using the canonical AR cycle (Staron, 2020) to structure and reflect on the process.
Our AR framing is particularly relevant in software engineering (SE), where controlled experiments, case studies, and surveys dominate, while AR remains comparatively rare (Petersen et al., 2014). Sjøberg et al. (Sjøberg et al., 2007) observe that part of the reason is a limited understanding of AR’s role in SE. By analyzing the AIBOM specification development effort through AR lens, we offer a concrete example of its application in the wild to the software engineering body of knowledge.
@@
Diagnosing Identifying a practical problem WG formation and scoping: Addressed the lack of a standardized way to describe AI components for traceability, compliance, and transparency.
Action Planning Designing an intervention Field definition and framework design: Extended SPDX to include AI artifacts by analyzing model cards, datasheets, and existing practices.
@@ 7. Lessons Learned and the Road Ahead
This section first distills the key lessons learned from our experience developing the AIBOM standard in an open, multi-stakeholder setting, and then outlines the road ahead by discussing how these lessons inform future extensions, particularly in the context of foundation model-powered software.
@@ 7.1. Lessons Learned
Present Work-in-Progress Early and Solicit Feedback Often. Public presentations of early drafts, in formal venues such as OSS Summit sessions (see Section 3.3) and other WGs provided critical feedback between mid-2022 and 2024. This early engagement not only strengthened the evolving standard but also built practitioner trust and buy-in. Feedback from outside our WG often prompted significant design changes. For instance, input from the OpenChain WG led to the introduction of a standardsCompliance field to capture which standards an AI model or dataset already adheres to, which is a key requirement for audits. A question during an OSS NA Vancouver panel about the queryability of fields such as energyConsumption and metrics prompted us to replace free-form text with structured types. Similarly, discussions at OSS Summit Japan 2022 revealed that tracking AI and data lineage separately would complicate adoption, leading us to merge them into a unified dataCollectionProcess field. Based on our experience, in line with previous studies (Centre., 2019), we suggest that continuous, feedback-driven development is not merely advantageous but essential for open standards in fast-moving domains like AI.
@@
Prioritize Adoption Over Comprehensiveness.  Early drafts that attempted to capture every conceivable detail of an AI system consistently faced pushback from practitioners. Most organisations simply do not maintain information at that level of granularity, and a standard that demands it becomes impractical (Jespersen, 1995; Isaak, 2005). A similar pattern is evident with model cards, which prescribe numerous fields, and as a consequence, over 60% of models and 70% of datasets on Hugging Face lack a complete model card (Stalnaker et al., 2025; Oreamuno et al., 2024). Several prior studies document similar resistance to heavyweight standards (Pfleeger et al., 2002; Schmidt, 2000).
@@
We therefore optimized our AIBOM specification for adoption by defining a small set of readily recordable required fields and enforcing strict entry criteria. In some cases, we intentionally excluded ambitious goals to improve practicality. For example, rather than attempting to enumerate every potential form of bias, we introduced a single knownBias field to capture only documented biases. Additional fields were deferred to future releases, enabling incremental evolution as practices mature.
@@
We also found that two factors: metadata scale and availability, strongly influence adoption. For instance, a comprehensive AIBOM for a self-driving car could span several gigabytes, making generation, storage, and maintenance impractical. At the same time, much of the desired metadata simply does not exist in accessible form. Even foundational datasets such as ImageNet and CIFAR-10 do not fully disclose their data sources (Rajbahadur et al., 2021). Supporting fields that can be automatically collected or derived are therefore essential. Recognizing this, we are now collaborating with the SPDX Implementers WG (SPDX Project / Linux Foundation, 2025) and CISA SBOM-o-RAMA (Cybersecurity and Infrastructure Security Agency (2025), CISA) to improve automation support in future iterations. These experiences highlight a broader reality: premature attempts at exhaustive coverage can undermine adoption. While similar lessons have been echoed several decades ago during POSIX standardization (Jespersen, 1995; Isaak, 2005), they ring true even today. Prioritizing a minimal, usable core builds early momentum and a foundation for richer requirements as the ecosystem matures.
@@
Plan for Churn and Conflict from Day One. In long-running, community-driven standardization efforts, both contributor turnover and conflicting priorities are inevitable, and planning for them early is essential. Several participants deeply involved in shaping the initial drafts disengaged as their organizational priorities shifted, while others joined later and made key contributions during specific phases. This irregular participation often left critical work unfinished, slowing progress. In response, we established a rotating core group with decision-making authority to ensure continuity while still enabling broader community contributions.
@@
Both subtle and overt disagreements were also a constant reality. In a multi-stakeholder setting with competing interests, participants frequently proposed adding, removing, or rewording fields, often with vague or narrowly scoped motivations. To prevent these debates from stalling progress, we consistently applied the evidence-based acceptance criteria defined in Section 7.1. This process significantly reduced friction and kept the WG focused on shared priorities rather than endless debate. As the AI ecosystem continues to diversify, sustaining continuity amid organizational churn and aligning competing priorities will become even more challenging. Embedding these mechanisms early helps ensure that open standards remain both stable and adaptable over time.

### v2

@@ 5.1. Regulatory and Standards Alignment
Methodology. We analysed the alignment between fields captured in the AIBOM specification and the regulatory and ethical requirements defined by the standards and regulations shown in Table 2. The analysis was performed by the fourth and fifth author, who have extensive experience in standards development and regulatory alignment, and was further independently reviewed by first and the sixth authors for additional validation.
EU AI Act., we examined the obligations defined in Articles 49 and 60, which govern mandatory registration in the EU high-risk AI database prior to market entry or testing in sandbox. Required information was grouped into four categories: Identification, System Details, Verification Details, and Application Details, comprising a total of 14 subcategories (e.g., provider contact, intended purpose, classification, testing plan, and involved parties).
US and EU medical device regulations, we analysed our AIBOM specification’s alignment with Article 10 of Regulation (EU) 2017/745 and FDA cybersecurity guidance. Information elements were organised into three categories: package details, model details, and data details.
@@
Results. Our analysis shows that the SPDX 3.0 AIBOM specification provides strong coverage of regulatory, safety, and ethical requirements. We found that we could successfully represent 13 of 14 information obligations under Articles 49 and 60 of the EU AI Act, capturing all required elements from US and EU medical device regulations. Our validation shows that AIBOM specification aligns with over 40 subclauses across eight IEEE 7000 series standards i.e., more than 90% overall coverage indicating practical readiness for compliance adoption.
@@ 5.2. Validation against Key Industry Use Cases
Motivation. While regulatory alignment ensures that our AIBOM specification meets legal and compliance obligations, its practical value ultimately depends on how well it addresses real-world needs identified by industry stakeholders. In June 2025, the SBOM for AI Tiger Team (SBOM for AI Tiger Team, 2025), convened by the U.S. Cybersecurity and Infrastructure Security Agency (CISA), articulated six foundational use cases that any AI-focused SBOM standard must support. These use cases span the full lifecycle of AI system deployment: regulatory compliance, Vulnerability and Incident Management, legal and intellectual property assurance, third-Party AI Risk Management, Open Source Model Risk Assessment and Model Lifecycle and Asset Management. A detailed overview of these usecases can be found here (for AI Tiger Team, 2025).
Validating our AIBOM specification against these use cases is therefore essential to demonstrate that the specification is not only standards-compliant but also operationally relevant, usable by organisations in practice, and capable of supporting the core security, legal, and governance workflows envisioned by the broader AIBOM consumer ecosystem.
Methodology. The CISA “SBOM for AI” use cases are intentionally high-level; however, assessing whether our AIBOM specification is capable of capturing the metadata outlined in them requires concrete, testable requirements. We therefore operationalised these use cases by consolidating them into broader themes, assembling a focused evidence base from recent studies, extracting atomic requirements, and mapping them to SPDX 3.0 AIBOM fields before performing external validation.
@@
Step 2: Build an evidence base per theme. For each theme, we identified relevant research papers and technical reports that analyse concrete instances of the use cases. For example, Rajbahadur et al. (Rajbahadur et al., 2021), which outlines methods for conducting dataset license compliance analysis, was mapped to the Compliance category. All selected studies are listed in Table 3.
@@
Step 5: Internal and external validation. The first author independently verified the extractions and mappings. We then presented the traceability matrix and representative examples to the AI SBOM Tiger Team (29 Sept 2025, attended by 12 participants) and the SPDX AI and Dataset WG (attended by four participants). Authors of this paper who were part of the working group did not participate in the validation exercise. Feedback was solicited on whether AIBOM could be effectively operationalised to address the six use cases. Meeting minutes from both working group sessions are publicly available. [^2] SBOM for AI (AIBOM) Tiger Team working group meeting minutes https://docs.google.com/document/d/1IpXG7XBOJnPl_hwFf3JZkDaFb0k2CnI0/edit?usp=sharing&ouid=110194678381965933391&rtpof=true&sd=true
Results. Our analysis demonstrated that the AIBOM specification could represent all 46 distinct requirements extracted from the eight studies reviewed across the three consolidated use-case categories. Table 3 summarises the number of AIBOM fields across the AI, Dataset, and other SPDX profiles that satisfied the requirements identified in each study. In several cases, a single requirement could be addressed by multiple fields, underscoring the flexibility and composability of our AIBOM specification. A detailed traceability matrix, including the full set of extracted requirements and their field-level mappings, is available in our extended technical report (Rashno, 2025).
@@
Open Source Risk Management (Thiel, 2023) 6 5 4
@@ 5.3. Validation based on Practitioner Interviews
Methodology. This validation occurred prior to the official AIBOM release, ensuring participants identified essential fields based on practical needs rather than reacting to the limitations of a finalised specification. Our methodology is detailed below.
Participant recruitment. We recruited ten practitioners via professional networks and targeted emails between 2022 and mid-2023. Participants were selected based on their expertise in SBOM development or AI software engineering to ensure a multi-disciplinary evaluation of the AIBOM framework.
Table 4 details the ten participants, who represent diverse work domains, experience levels, and potential future roles within the AIBOM ecosystem.
@@
Interview protocol. After refining an initial questionnaire based on feedback from two academic and industry experts, we finalised an 18-question protocol. Interviews explored four themes: (1) participants’ experience in AI and software development; (2) working definitions of AI software/system traceability; (3) current traceability practices and tools; and (4) expectations for the new AIBOM specification.
We conducted semi-structured video interviews (30–60 minutes) to allow for improvisation and exploration (Wohlin et al., 2012). To avoid bias, the proposed AIBOM fields were not disclosed during the sessions. With participant consent, interviews were recorded, transcribed via software, and manually verified for accuracy.
Transcript coding and analysis. Two authors initially co-coded two transcripts to induce a baseline coding scheme. The remaining transcripts were then coded independently. Following this, the authors met to filter irrelevant codes and reach a consensus on categories. Finally, we computed the overlap between these participant-derived concepts and our proposed AIBOM specification fields.
@@
Results. Excluding Configuration and Deployment fields—reserved for future versions due to capture complexity—all participant-proposed fields map to the SPDX AIBOM 3.0 specification. Table 5 details this mapping between participant suggestions and the specification.
The primary theme emerging from the interviews was the widespread inadequacy of current traceability practices for AI systems, which participants described as ad hoc and insufficient for ensuring proper governance. This sentiment was powerfully articulated by an executive director (P2), who stated, ”I think it’s a general mystery across the industry… I don’t think that [traceability] adequately exists within the AI community around algorithms and training sets to be able to demonstrate that traceability from an auditing standpoint.” This lack of a standardised, auditable record is exacerbated by real-world development pressures. A machine learning engineer (P7) from a startup described this vividly: ”It’s something that we lack and it’s because… the general nature of a startup, we usually deprioritise documentation and it usually also ends with a lot of pain for us… because we don’t have the documentation, I usually spend so much time trying to explain anyway.”
Beyond identifying these gaps, participants also provided clear insights into effective traceability requirements. When asked what information they need when consuming a pre-trained model, a lead data scientist (P1) offered a detailed wishlist: ”What do I look for? I look for license… support… When, which training data was used, what demographic was used, and what biases do they have? …And these are stuff that you want captured in the document, because that’s what I’m looking for [as a model consumer]… what’s the [reported] accuracy? How did you test it?” P1’s requirements—spanning licensing, provenance, and evaluation—directly align with the AIBOM framework’s core categories, confirming that its design addresses practitioner needs and will guide the specification’s evolution.
@@ 5.4. Field Validation in Industrial Setting
We evaluated AIBOM’s practical suitability through an external study at a large multinational software organisation. From June to September 2024, a multi-disciplinary team of security architects, data scientists, DevOps experts, and AI developers mapped the overlap between our specification fields and those required during the organisation’s AI system development.
Methodology. The organisation already had a checklist for developers to evaluate how their AI software development affects the internal ethics policy. Moreover, the legal department had a checklist to assess the legal impact of each AI-related software release. The researchers investigated how many of the fields in these checklists could be populated if the organisations already had AIBOMs constructed as part of the AI software development process. Furthermore, they investigated whether our AIBOM specification can be used to automatically generate model cards. Then, they investigated to what extent AIBOM generation for third-party models can be automated using web scraping or external APIs, such as Hugging Face (Face, 2025).
Results. Validation revealed that AIBOM fields cover all internal checklist requirements for developers, proving its practical comprehensiveness. They observed that 60% of existing model card fields are extractable from the fields in the AIBOM specification. Furthermore, external APIs and web scraping successfully populated 40% of fields for third-party models, demonstrating that AIBOM generation can be partially automated to reduce manual effort.
@@ 6. Standards Development as Action Research
A primary contribution of this work is demonstrating how the development of a complex, multi-stakeholder standard can be systematically understood through the lens of Action Research (AR). By retrospectively framing our effort to extend SPDX and create an AIBOM specification as an AR process, we illustrate how community-driven standards development can maintain a disciplined, research-grounded approach within rapidly evolving domains.
While not initially conceived as AR, this project’s conduct and outcomes align closely with its core principles. AR focuses on changing practice through real-world intervention (Baskerville, 1999; Elden and Chisholm, 1993; Ferrario et al., 2014); similarly, our objective was to develop a functional, community-driven standard for AI transparency and compliance. We therefore adopt the canonical AR cycle (Staron, 2020) to retrospectively structure and reflect on this process.
Our AR framing is particularly relevant in software engineering (SE), where controlled experiments, case studies, and surveys dominate, while AR remains comparatively rare (Petersen et al., 2014). Sjøberg et al. (Sjøberg et al., 2007) attribute this rarity partly to a limited understanding of AR’s role in SE. By analysing the AIBOM specification development via an AR lens, we contribute a concrete example of its application in the wild to the software engineering body of knowledge.
@@
Diagnosing Identifying a practical problem WG formation and scoping: Addressed the lack of a standardised way to describe AI components for traceability, compliance, and transparency.
Action Planning Designing an intervention Field definition and framework design: Extended SPDX to include AI artefacts by analysing model cards, datasheets, and existing practices.
@@ 7. Lessons Learned and the Road Ahead
This section distills lessons from developing the AIBOM standard in a multi-stakeholder setting and discusses how these insights inform future extensions for foundation model-powered software.
@@ 7.1. Lessons Learned
Present work-in-progress early and solicit feedback often. Public presentations of early drafts, in venues such as Open Source Summit (OSS) sessions (see Section 3.3) and other WGs provided critical feedback between mid-2022 and 2024. This early engagement not only strengthened the evolving standard but also built practitioner trust and buy-in. Feedback from outside our WG often prompted significant design changes. For instance, input from the OpenChain WG led to the introduction of a standardsCompliance field to capture standards an AI model or dataset already adheres to, which is a key requirement for audits. A question during an OSS North America 2023 panel about the queryability of fields such as energyConsumption and metrics prompted us to replace free-form text with structured types. Similarly, discussions at OSS Japan 2022 revealed that tracking AI and data lineage separately would complicate adoption, leading us to merge them into a unified dataCollectionProcess field. Based on our experience, in line with previous studies (Centre., 2019), we suggest that continuous feedback-driven development is not only advantageous but essential for open standards in fast-moving domains like AI.
@@
Prioritise adoption over comprehensiveness.  Early drafts that attempted to capture every conceivable detail of an AI system consistently faced pushback from practitioners. Most organisations simply do not maintain information at that level of granularity, and a standard that demands it becomes impractical (Jespersen, 1995; Isaak, 2005). A similar pattern is evident with model cards, which prescribe numerous fields; over 60% of models and 70% of datasets on Hugging Face lack a complete model card (Stalnaker et al., 2025; Oreamuno et al., 2024). Several prior studies document similar resistance to heavyweight standards (Pfleeger et al., 2002; Schmidt, 2000). Another responsible AI field study reports an organisation expressing “a real fear that they may adopt principles that they cannot uphold in practice” (Ruster and Davis, 2025).
@@
We therefore optimised our AIBOM specification for adoption by defining a small set of readily recordable required fields and enforcing strict entry criteria. In some cases, we intentionally excluded ambitious goals to improve practicality. For example, rather than attempting to enumerate every potential form of bias, we introduced a single knownBias field to capture only documented biases. Additional fields were deferred to future releases, enabling incremental evolution as practices mature.
@@
We also found that two factors: metadata scale and availability, strongly influence adoption. For instance, a comprehensive AIBOM for a self-driving car could span several gigabytes, making generation, storage, and maintenance impractical. At the same time, much of the desired metadata simply does not exist in accessible form. Even foundational datasets such as ImageNet and CIFAR-10 do not fully disclose their data sources (Rajbahadur et al., 2021). Supporting fields that can be automatically collected or derived are therefore essential. Recognising this, we are now collaborating with the SPDX Implementers WG (SPDX Project / Linux Foundation, 2025) and CISA SBOM-o-RAMA (Cybersecurity and Infrastructure Security Agency, 2025b) to improve automation support in future iterations. These experiences highlight a broader reality: premature attempts at exhaustive coverage can undermine adoption. While similar lessons have been echoed several decades ago during POSIX standardisation (Jespersen, 1995; Isaak, 2005), they ring true even today. Prioritising a minimal, usable core builds early momentum and a foundation for richer requirements as the ecosystem matures.
@@
Plan for churn and conflict from day one. In long-running, community-driven standardisation, contributor turnover and conflicting priorities are inevitable; planning for them early is essential. Several participants disengaged as organisational priorities shifted, while others joined late to contribute to specific phases. This irregular participation often left critical tasks unfinished, hindering progress. Consequently, we established a rotating core group with decision-making authority to ensure continuity alongside broader community input.
@@
Disagreements were a constant reality in this multi-stakeholder setting. Participants frequently proposed modifications with vague or narrow motivations, risking developmental stalls. To mitigate this, we applied the evidence-based acceptance criteria defined in Section 7.1, significantly reducing friction and maintaining focus on shared priorities. As the AI ecosystem diversifies, sustaining continuity amid organisational churn and aligning competing interests will grow more challenging. Embedding these governance mechanisms early ensures open standards remain stable yet adaptable.